OpenStudy (maddiie):

Use the perfect square trinomial formuals to factor. Be sure to factor out any common factors. 16y^2-8y+1

OpenStudy (amistre64):

well; the name of it seems rather author specific...

OpenStudy (amistre64):

divide out the 16; and then complete the square

OpenStudy (anonymous):

we have 16 = 4 * 4 and -4-4 = -8 so the factors are (4y - 1)(4y-1) or (4y-1)^2

OpenStudy (amistre64):

16y^2-8y+1 ----------- 16 y^2 -(1/2)y + (1/16) ; (1/2)/2 = 1/4, sqaure it y^2 -(1/2)y +1/16 - 1/16 + 1/16 ; group it (y^2 -(1/2)y +1/16) +(-1/16 + 1/16) (y-1/4)^2 +0

OpenStudy (amistre64):

16(y-(1/4))^2 perhaps ╠╬╩